# Mixed Diffusion for 3D Indoor Scene Synthesis

### v1

@@ Abstract
Realistic conditional 3D scene synthesis significantly enhances and accelerates the creation of virtual environments, which can also provide extensive training data for computer vision and robotics research among other applications. Diffusion models have shown great performance in related applications, e.g., making precise arrangements of unordered sets. However, these models have not been fully explored in floor-conditioned scene synthesis problems. We present MiDiffusion, a novel mixed discrete-continuous diffusion model architecture, designed to synthesize plausible 3D indoor scenes from given room types, floor plans, and potentially pre-existing objects. We represent a scene layout by a 2D floor plan and a set of objects, each defined by its category, location, size, and orientation. Our approach uniquely implements structured corruption across the mixed discrete semantic and continuous geometric domains, resulting in a better conditioned problem for the reverse denoising step. We evaluate our approach on the 3D-FRONT dataset. Our experimental results demonstrate that MiDiffusion substantially outperforms state-of-the-art autoregressive and diffusion models in floor-conditioned 3D scene synthesis. In addition, our models can handle partial object constraints via a corruption-and-masking strategy without task specific training. We show MiDiffusion maintains clear advantages over existing approaches in scene completion and furniture arrangement experiments.
@@ 1 Introduction
We focus on object-based 3D indoor scene synthesis in which the goal is to predict a set of objects characterized by semantic labels and geometric arrangements. Results using state-of-the-art autoregressive models show that, while it is relatively easy to predict reasonable object labels given a room type, the main bottleneck is generating accurate geometric arrangements of objects, i.e., predicting realistic positions and orientations. Autoregressive models might fail because they cannot revert a bad prediction in an earlier previous step. Therefore, “de-noising” approaches that iteratively refine predicted results are better suited for this type of problems [46].
In this work, we propose to adopt diffusion models [35] for this problem, as the iterative denoising scheme is well suited for improving object geometry in 3D scenes. Denoising Diffusion Probabilistic Models (DDPM) [15] proposed Gaussian corruption and denoising of continuous domain data, resulting in efficient training of diffusion models for image synthesis tasks. Discrete Denoising Diffusion Probabilistic Models (D3PM) [2] extended the pioneering work of Hoogeboom et al. [16] to structured categorical corruption processes for data in discrete state space. Unlike continuous diffusion models, discrete diffusion involves data corruption and recovery using transition probabilities between discrete states. Since object attributes are naturally defined on both discrete and continuous domains, we propose to combine these approaches to form a mixed discrete-continuous diffusion model for 3D scene synthesis problems.
There are a few recent works adopting diffusion models or iterative denoising strategies for 3D scene synthesis problems [17, 13, 46, 41]. The closest work to ours is DiffuScene [41], which generates 3D indoor scenes without floor plan constraints using DDPM by first converting semantic labels to one-hot encoding vectors in the continuous domain and then using argmax to retrieve label predictions. We instead focus on floor-conditioned problems, as these are more practical setups in automatic design of real or virtual worlds. We present three technical contributions in this paper:
@@
  • We design a time variant transformer-based denoising network with floor plan features as condition vectors for our mixed diffusion approach.
@@
We compare MiDiffusion against state-of-the-art autoregressive and diffusion models in floor-conditioned 3D scene synthesis problems. Experimental evaluation using the common 3D-FRONT [9] dataset demonstrates that our approach generates more realistic scene layouts, outperforming existing methods. We provide ablation studies supporting our design of the mixed diffusion model and the denoising network. Without re-training, we also show experimental evaluations on scene completion and furniture arrangement, confirming that our approach maintains superior performance on larger scenes. Code and data are available at https://github.com/MIT-SPARK/MiDiffusion.
@@ 2 Related Works
3D Scene Synthesis. 3D scene synthesis is typically achieved by generating a set of object layouts from scratch. Classical approaches in the computer graphics community typically employ procedural modeling to apply a set of functions that capture object relationships in indoor [50, 25, 49, 32] and outdoor [40, 25, 49, 29, 18, 5] settings. A handful of methods are based on graph representations of the scene. Meta-Sim [18] and Meta-Sim2 [5] learn to modify attributes of scene graphs obtained from probabilistic scene grammars to match synthesized and target images. SG-VAE [30] learns a grammar-based auto-encoder to capture relationships and decode the latent space to a parse tree. Most recently, it is common to learn inter-object relationships implicitly without specifying hand-crafted rules. Existing models include feed-forward networks [53], VAEs [30, 47], GANs [48], and autoregressive models [32, 33, 43, 23, 53, 54, 27, 44, 28]. The state-of-the-art autoregressive models [44, 28] are built with a transformer [42] backbone, capable of well modelling object interactions.
@@
Diffusion Models. Diffusion probabilistic models [38, 15, 2, 26, 36] are generative models defined by two Markov processes. The forward process slowly injects random noise to the data, whereas the reverse denoising process recovers the data. Conceptually, this approach has connections with denoising score matching methods [37, 39]. Following the pioneering work by [35], diffusion models first gained popularity in 2D image synthesis [19, 36], outperforming existing works [20, 6]. Starting from Denoising Diffusion Probabilistic Models (DDPM), diffusion models typically work with latent variables in the continuous domain. However, it is more natural to represent discrete variables, such as text, in discrete state space. Discrete diffusion models were first applied to text generation in argmax flow [16], and then Diffusion Probabilistic Models (D3PMs) [2] and VQ-Diffusion [11] showed strong results on character level image generation and text-to-image synthesis.
@@
Diffusion Models for Layout Synthesis. Very recently, diffusion models became popular for layout synthesis, including document layout generation [17, 13], 3D scene synthesis [41], furniture re-arrangement [46, 41], graph-conditioned 3D layout generation [52, 24], and text conditioned scene synthesis [7]. LayoutDM [17] applies diffusion models in discrete state space. After discretizing position and size to a fixed number of bins, they use the discrete corruption process by VQ-Diffusion [11] to train the reverse transformer network for predicting category, position, and size for document layout generation. Most related to ours, DiffuScene [41] predicts 3D scene layout using DDPM after converting semantic labels to one-hot encodings to allow joint diffusion with geometric attributes in the continuous domain. In contrast, our diffusion model works directly on the mixed discrete-continuous domains. Moreover, we design a novel corruption-and-masking strategy for reverse diffusion under partial object constraints, without the need to re-train any models.
@@ 3 Preliminary: Diffusion Models
$\begin{split}L_{vb}=\mathbb{E}_{q({\bm{x}}_{0})}[&\underbrace{D_{\mathrm{KL}}\left(q({\bm{x}}_{T}|{\bm{x}}_{0})||p({\bm{x}}_{T})\right)}_{L_{T}}+\sum_{t=2}^{T}\underbrace{{\mathbb{E}}_{q({\bm{x}}_{t}|{\bm{x}}_{0})}[D_{\mathrm{KL}}(q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})||p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t}))]}_{L_{t-1}}\\ &+\underbrace{{\mathbb{E}}_{q({\bm{x}}_{1}|{\bm{x}}_{0})}[-\log p_{\theta}({\bm{x}}_{0}|{\bm{x}}_{1})]}_{L_{0}}].\end{split}$ (2)
@@
We present MiDiffusion, a mixed discrete-continuous diffusion model for 3D indoor scene synthesis. We assume each scene ${\cal S}$ is in a world frame with the origin at the center. It consists of a floor plan and at most $N$ objects. We denote the floor plan by a conditional vector ${\bm{y}}$. Each object in the scene is characterized by its semantic label $z\in\{1,2,\dots,C\}$, centroid position ${\bm{t}}\in{{\mathbb{R}}^{3}}$, bounding box size ${\bm{s}}\in{{\mathbb{R}}^{3}}$, and rotation angle around the vertical axis $\phi\in SO(2)$. Inspired by LEGO-Net [46], we represent $\phi$ as $[\cos(\gamma),\sin(\gamma)]^{\mathsf{T}}$ to maintain continuity on $SO(2)$ [55]. We concatenate all geometric attributes to ${\bm{x}}=({\bm{t}},{\bm{s}},\phi)$. Then, we can represent an 3D indoor scene as ${\cal S}=(\{(z^{i},{\bm{x}}^{i})\}_{1\leq i\leq N},{\bm{y}}$). We designate the last semantic label $z=C$ to an “empty” label to work with scenes containing less than $N$ objects and define geometric attributes to all zeros for an “empty” object.
The problem of synthesising 3D scenes can be viewed as learning a network $\theta$ that maximizes $\log p_{\theta}(z_{0}^{i},{\bm{x}}_{0}^{i}|{\bm{y}})$ over all objects and across all scenes, i.e., the same variational bound in Eq. (2) but over two domains. We drop the superscript $i$ for simplicity.
@@ 4.1 Diffusion in Mixed Discrete-Continuous Domains
In the backward diffusion process, we can design a network $\theta$ that computes probability distributions over latent variables by domains to be consistent with the forward process:
$p_{\theta}\left(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}\right):={\tilde{p}}_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}})\cdot{\hat{p}}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}}).$ (12)
@@
$\begin{split}L_{t-1}^{mixed}=&{\mathbb{E}}_{q\left(z_{t},{\bm{x}}_{t}|z_{0},{\bm{x}}_{0}\right)}[D_{\mathrm{KL}}(q\left(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},z_{0},{\bm{x}}_{0}\right)||p_{\theta}\left(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}\right))]\\ =&\underbrace{{\mathbb{E}}_{\tilde{q}(z_{t}|z_{0})}[D_{\mathrm{KL}}(\tilde{q}(z_{t-1}|z_{t},z_{0})||\tilde{p}_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}))]}_{L_{t-1}^{D3PM}}\\ &+\underbrace{{\mathbb{E}}_{\hat{q}({\bm{x}}_{t}|{\bm{x}}_{0})}[D_{\mathrm{KL}}(\hat{q}({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})||\hat{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}}))]}_{L_{t-1}^{DDPM}},\end{split}$ (13)
$\begin{split}L_{0}^{mixed}=&{\mathbb{E}}_{q\left(z_{1},{\bm{x}}_{1}|z_{0},{\bm{x}}_{0}\right)}\left[-\log p_{\theta}\left(z_{0},{\bm{x}}_{0}|z_{1},{\bm{x}}_{1},{\bm{y}}\right)\right]\\ =&\underbrace{{\mathbb{E}}_{\tilde{q}\left(z_{1}|z_{0}\right)}\left[-\log\tilde{p}_{\theta}\left(z_{0}|z_{1},{\bm{x}}_{1},{\bm{y}}\right)\right]}_{L_{0}^{D3PM}}+\underbrace{{\mathbb{E}}_{\hat{q}\left({\bm{x}}_{1}|{\bm{x}}_{0}\right)}\left[-\log\hat{p}_{\theta}\left({\bm{x}}_{0}|{\bm{x}}_{1},z_{1},{\bm{y}}\right)\right]}_{L_{0}^{DDPM}}.\end{split}$ (14)
@@
Combining Eq. (13), (14), we get the exact sum of variational bounds in D3PM and DDPM. Intuitively, summing over discrete and continuous losses is similar to summing losses over each coordinate given multi-dimensional data in vanilla DDPM or D3PM. We use a simplified version of $L_{vb}^{DDPM}$ after re-parameterizing the forward process as suggested in DPPM to improve training stability and efficiency. We add an auxiliary loss, $L_{aux}^{D3PM}$, as proposed in D3PM to encourage good predictions for $z_{0}$ at each time step. We include re-parameterization steps as suggested by DDPM and D3PM in Appendix A. The combined loss for MiDiffusion is:

### v2

@@ Abstract
Generating realistic 3D scenes is an area of growing interest in computer vision and robotics. However, creating high-quality, diverse synthetic 3D content often requires expert intervention, making it costly and complex. Recently, efforts to automate this process with learning techniques, particularly diffusion models, have shown significant improvements in tasks like furniture rearrangement. However, applying diffusion models to floor-conditioned indoor scene synthesis remains under-explored. This task is especially challenging as it requires arranging objects in continuous space while selecting from discrete object categories, posing unique difficulties for conventional diffusion methods. To bridge this gap, we present MiDiffusion, a novel mixed discrete-continuous diffusion model designed to synthesize plausible 3D indoor scenes given a floor plan and pre-arranged objects. We represent a scene layout by a 2D floor plan and a set of objects, each defined by category, location, size, and orientation. Our approach uniquely applies structured corruption across mixed discrete semantic and continuous geometric domains, resulting in a better-conditioned problem for denoising. Evaluated on the 3D-FRONT dataset, MiDiffusion outperforms state-of-the-art autoregressive and diffusion models in floor-conditioned 3D scene synthesis. Additionally, it effectively handles partial object constraints via a corruption-and-masking strategy without task-specific training, demonstrating advantages in scene completion and furniture arrangement tasks.
@@ 1 Introduction
To assist these applications, we focus on object-based 3D indoor scene synthesis, which aims at predicting a set of objects with both semantic labels and consistent geometric arrangements. These predicted objects can be rendered into photorealistic scenes using furniture CAD models or other downstream shape prediction methods [39]. State-of-the-art works have recently shown that it is relatively easy to predict reasonable object labels, yet, generating accurate geometric arrangements of objects, i.e. predicting realistic positions and orientations, is still very challenging[32, 49]. In particular, as most of the existing approaches are based on autoregressive models, these models tend to quickly start diverging upon a bad prediction in an earlier step. Therefore, “de-noising” approaches that jointly refine all objects are thus better suited for this type of problems [51].
We propose to adopt diffusion models [40] for this problem, as the iterative denoising scheme is well suited for improving object geometry in 3D scenes. Denoising Diffusion Probabilistic Models (DDPM) [17] adopt Gaussian corruption and denoising of continuous domain data, resulting in efficient training of diffusion models for image synthesis tasks. Discrete Denoising Diffusion Probabilistic Models (D3PM) [2] extended the pioneering work of Hoogeboom et al. [19] to structured categorical corruption processes for data in discrete state space. Unlike continuous diffusion models, discrete diffusion involves data corruption and recovery using transition probabilities between discrete states. Since object attributes are naturally defined on both discrete and continuous domains, we derive a novel mixed discrete-continuous diffusion model, optimally suited for 3D scene synthesis problems.
A handful of works have recently also adopted diffusion models or iterative denoising strategies for 3D scene synthesis problems [20, 15, 51, 46]. The closest work to ours is DiffuScene [46], which generates 3D indoor scenes using DDPM by first converting semantic labels to one-hot encoding vectors and then combining with geometric features in the continuous domain. While this works overall well, for categorical data, discrete diffusion captures dependencies more naturally and more effectively. Therefore, we propose a novel mixed discrete-continuous denoising formulation which which jointly corruptss and denoises discrete object categories and continuous geometric arrangements, leading to superior results. Note that we focus on floor-conditioned problems, as these are more practical setups in automatic design of real or virtual worlds. We present three technical contributions:
@@
  • We propose a modified version of a time variant transformer denoising network that is compatible with our mixed diffusion formulation and floor plan conditioning.
@@
We compare MiDiffusion against state-of-the-art autoregressive and diffusion models on the common 3D-FRONT [10] dataset, demonstrating that our approach generates more realistic scene layouts, outperforming all existing methods. Code and data are available at https://github.com/MIT-SPARK/MiDiffusion.
@@ 2 Related Works
In this section, we review key works in 3D scene synthesis, followed by an overview of diffusion models, and conclude with relevant studies on layout synthesis using diffusion.
@@
3D Scene Synthesis. 3D scene synthesis is typically achieved by generating a set of object layouts from scratch. Classical approaches in the computer graphics community typically employ procedural modeling to apply a set of functions that capture object relationships in indoor [56, 29, 55, 36] and outdoor [45, 29, 55, 33, 21, 5] settings. A handful of methods are based on graph representations of the scene. Meta-Sim [21] and Meta-Sim2 [5] learn to modify attributes of scene graphs obtained from probabilistic scene grammars to match synthesized and target images. SG-VAE [34] learns a grammar-based auto-encoder to capture relationships and decode the latent space to a parse tree. Most recently, it is common to learn inter-object relationships implicitly without specifying hand-crafted rules. Existing models include feed-forward networks [60], VAEs [34, 53], GANs [54], and autoregressive models [36, 37, 48, 26, 60, 62, 31, 49, 32]. The state-of-the-art autoregressive models [49, 32] are commonly built with a transformer [47] backbone, making them capable of modeling object interactions through attention.
@@
Diffusion Models. Diffusion probabilistic models [43, 17, 2, 30, 41] are generative models defined by two Markov processes. The forward process slowly injects random noise to the data, whereas the reverse denoising process recovers the data. Conceptually, this approach has connections with denoising score matching methods [42, 44]. Following the pioneering work by [40], diffusion models first gained popularity in 2D image synthesis [23, 41], outperforming existing works [22, 6]. Starting from Denoising Diffusion Probabilistic Models (DDPM) [17, 40], diffusion models typically work with latent variables in the continuous domain. However, it is more natural to represent discrete variables, such as text, in discrete state space. Discrete diffusion models were first applied to text generation in argmax flow [19]. In the following, Diffusion Probabilistic Models (D3PMs) [2] and VQ-Diffusion [13] have further demonstrated strong results also on character level image generation and text-to-image synthesis.
@@
Diffusion Models for Layout Synthesis. Very recently, diffusion models became popular for layout synthesis, including document layout generation [20, 15], 3D scene synthesis [46], furniture re-arrangement [51, 46], graph-conditioned 3D layout generation [58, 59], text conditioned scene synthesis [27, 9, 12], and 3D scene mesh generation [52, 28]. LayoutDM [20] applies diffusion models in discrete state space. After discretizing position and size to a fixed number of bins, it uses the discrete corruption process by VQ-Diffusion [13] to train the reverse transformer network to predict category, position, and size for document layout generation. Most related to ours, DiffuScene [46] predicts 3D scene layouts using DDPM after converting semantic labels to one-hot encodings, allowing joint diffusion with geometric attributes in the continuous domain. In contrast, our model operates on mixed discrete-continuous domains, leading to more realistic scene layouts and increased robustness. Note that DiffuScene includes a simple demonstration of text-conditioned scene synthesis with a pretrained text parser. However, since we focus on joint semantic and geometric prediction, this line of work [7, 27, 39, 18] is not directly comparable to ours.
@@ 3 Preliminary: Diffusion Models
$\begin{split}L_{vb}=&\mathbb{E}_{q({\bm{x}}_{0})}[\underbrace{D_{\mathrm{KL}}\left(q({\bm{x}}_{T}|{\bm{x}}_{0})||p({\bm{x}}_{T})\right)}_{L_{T}}\\ &+\sum_{t=2}^{T}\underbrace{{\mathbb{E}}_{q({\bm{x}}_{t}|{\bm{x}}_{0})}[D_{\mathrm{KL}}(q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})||p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t}))]}_{L_{t-1}}\\ &+\underbrace{{\mathbb{E}}_{q({\bm{x}}_{1}|{\bm{x}}_{0})}[-\log p_{\theta}({\bm{x}}_{0}|{\bm{x}}_{1})]}_{L_{0}}].\end{split}$ (2)
@@
Figure 1: Denoising network. The time-variant decoder backbone takes object features as input, conditioned on the floor plan feature. We utilize the time-variant transformer decoder block from VQ-Diffusion [13].
@@
We present MiDiffusion, a mixed discrete-continuous diffusion model for 3D indoor scene synthesis. We assume each scene ${\cal S}$ is in a world frame with the origin at the center. It consists of a floor plan and at most $N$ objects. We denote the floor plan by a conditional vector ${\bm{y}}$. Each object in the scene is characterized by its semantic label $z\in\{1,2,\dots,C\}$, centroid position ${\bm{t}}\in{{\mathbb{R}}^{3}}$, bounding box size ${\bm{s}}\in{{\mathbb{R}}^{3}}$, and rotation angle around the vertical axis $\phi\in SO(2)$. Inspired by LEGO-Net [51], we represent $\phi$ as $[\cos(\gamma),\sin(\gamma)]^{\mathsf{T}}$ to maintain continuity on $SO(2)$ [61]. We concatenate all geometric attributes to ${\bm{x}}=({\bm{t}},{\bm{s}},\phi)$. Then, we can represent a 3D indoor scene as ${\cal S}=(\{(z^{i},{\bm{x}}^{i})\}_{1\leq i\leq N},{\bm{y}}$). We designate the last semantic label $z=C$ to be an “empty” label to work with scenes containing less than $N$ objects and define geometric attributes to all zeros for an “empty” object.
The problem of synthesising 3D scenes can be viewed as learning a network $\theta$ that maximizes $\log p_{\theta}(z_{0}^{i},{\bm{x}}_{0}^{i}|{\bm{y}})$ over all objects and across all scenes, i.e. the same variational bound in Eq. (2) but over two domains. We drop the superscript $i$ for simplicity.
@@ 4.1 Diffusion in Mixed Discrete-Continuous Domains
In the backward diffusion process, we force the network $\theta$ to compute updates by domains conditioned on the previous latent variables:
$\begin{split}p_{\theta}&\left(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}\right)\\ &:={\tilde{p}}_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}})\cdot{\hat{p}}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}}).\end{split}$ (13)
@@
$\begin{split}&L_{t-1}^{mixed}={\mathbb{E}}_{q\left(z_{t},{\bm{x}}_{t}|z_{0},{\bm{x}}_{0}\right)}[D_{\mathrm{KL}}(q\left(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},z_{0},{\bm{x}}_{0}\right)\\ &\hskip 128.0374pt||p_{\theta}\left(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}\right))]\\ &=\underbrace{{\mathbb{E}}_{\tilde{q}(z_{t}|z_{0})}[D_{\mathrm{KL}}(\tilde{q}(z_{t-1}|z_{t},z_{0})||\tilde{p}_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}))]}_{L_{t-1}^{D3PM}}\\ &\quad+\underbrace{{\mathbb{E}}_{\hat{q}({\bm{x}}_{t}|{\bm{x}}_{0})}[D_{\mathrm{KL}}(\hat{q}({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})||\hat{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}}))]}_{L_{t-1}^{DDPM}},\end{split}$ (14)
$\begin{split}&L_{0}^{mixed}={\mathbb{E}}_{q\left(z_{1},{\bm{x}}_{1}|z_{0},{\bm{x}}_{0}\right)}\left[-\log p_{\theta}\left(z_{0},{\bm{x}}_{0}|z_{1},{\bm{x}}_{1},{\bm{y}}\right)\right]\\ &=\underbrace{{\mathbb{E}}_{\tilde{q}\left(z_{1}|z_{0}\right)}\left[-\log\tilde{p}_{\theta}\left(z_{0}|z_{1},{\bm{x}}_{1},{\bm{y}}\right)\right]}_{L_{0}^{D3PM}}\\ &\quad+\underbrace{{\mathbb{E}}_{\hat{q}\left({\bm{x}}_{1}|{\bm{x}}_{0}\right)}\left[-\log\hat{p}_{\theta}\left({\bm{x}}_{0}|{\bm{x}}_{1},z_{1},{\bm{y}}\right)\right]}_{L_{0}^{DDPM}}.\end{split}$ (15)
@@
Note that the resulting loss, i.e. Eq. (14) and (15), is the exact sum of variational bounds in D3PM and DDPM. The only constraint is that the denoising network has to obey Eq. (13). Intuitively, summing over discrete and continuous losses is similar to summing losses over each coordinate given multi-dimensional data in vanilla DDPM or D3PM. We use a simplified version of $L_{vb}^{DDPM}$ after re-parameterizing the forward process as suggested in DPPM to improve training stability and efficiency. We add an auxiliary loss, $L_{aux}^{D3PM}$, as proposed in D3PM to encourage good predictions for $z_{0}$ at each time step. We include re-parameterization steps as suggested by DDPM and D3PM in Appendix A. The combined loss for MiDiffusion is: